OpenStudy (anonymous):

Factor completely 5a2 + b. Prime a(5a + b) b(5a2) ab(5a + 1) @iGreen. @Directrix

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

A

Directrix (directrix):

Why do you think it is prime? @knightmare6 Just wondering.

OpenStudy (anonymous):

Just because one sec ill show you my work

OpenStudy (anonymous):

a(5a + b) = 5a^2 + ab b(5a^2) = 5a^2b b(5a+ 1) = 5a^2b + ab

Directrix (directrix):

Good work. I agree that the expression is prime. Although, I almost chose this: a(5a + b)

OpenStudy (anonymous):

Okay why